OpenStudy (anonymous):

Find the vertex of the following function. f(x) = -2x2 - 4x find the order pairs?

OpenStudy (anonymous):

oh this is easy: set y = o and solve for x, you will get two solutions for x then substitute one solution to the original equation and solve for y do that again with the other x value and you have two coordinates

OpenStudy (anonymous):

f(x) = -2x^2 -4x set f(x) = 0 0 = -2x^2 - 4x 0 = -2x( x + 2) -2x = 0 or x + 2 = 0 x = 0 or x = -2 Substitute the first solution into the original equation f(x) = -2x^2 - 4x f(0) = -2(0)^2 -4(0) f(0) = 0 * you got a y value for your first coordinate Substitute the second solution into the original equation f(x) = -2x^2 - 4x f(-2) = -2(-2)^2 - 4(-2) f(-2) = -2(4) + 8 f(-2) = -8 + 8 f(-2) = 0

OpenStudy (anonymous):

good job dada

OpenStudy (anonymous):

thanks oh i forgot to tell you the coordinates are : (0,0) and (-2,0)